1. A diesel engine has a compression ratio of 18 and cut- off takes place at 5% the stroke . calculate the air standard efficiency. Y=1.4
1
Expert's answer
2021-05-31T06:08:44-0400
"V_s=v_1-v_2=17v_2\\v_3=0.05(17v_2)+v_2=1.85v_2"

"\eta=1-\frac{1}{18^{0.4}}\frac{1.85^{1.4}-1}{1.4(1.85-1)}=0.639"